# Characterizing Agent-Based Model Dynamics via $\varepsilon$-Machines and Kolmogorov-Style Complexity

### v2

@@ 5 Case Study Outline: Caregiver–Elder Dyads
From a methodological standpoint, this demonstrates why symbolization cannot be regarded as a neutral preprocessing step. The quantile scheme was adequate for efforts, whose variability is driven primarily by cross-sectional attributes (income and elder ability) rather than by temporal feedback. In that case, the absence of causal states corresponds to genuine statistical independence. Yet for walkability and hours not cared, the same discretization distorted the data-generating process: it produced $\varepsilon$-machines that appeared memoryless despite substantive reasons to expect autocorrelation—spatial continuity in the former and dependency on caregiving load in the latter. In both cases, the failure lies not in the reconstruction algorithm but in the representation of the signal.
@@ 7 Discussion, Limitations, and Conclusions
Compared with prior computational-mechanics studies that focus on theoretical or synthetic sources, the present framework operationalizes intrinsic computation in an applied, data-driven ABM setting, enabling systematic evaluation of informational dynamics in social simulations.
Methodologically, this study bridges theoretical computational mechanics and applied complexity analysis by embedding $\varepsilon$-machine reconstruction within an agent-based modeling workflow. The proposed methodology integrates symbolic dynamics, causal-state reconstruction, and algorithmic compressibility to evaluate informational organization in agent-based simulations. This multi-level representation—from symbol transitions to $\varepsilon$-machines and compression metrics—yields a reproducible and quantitative description of emergent predictability and structural heterogeneity in simulated social systems.
Whereas intrinsic computation traditionally quantifies how a stationary process stores and transmits information, our approach treats each dyad-level time series as a locally adaptive source whose causal states evolve under behavioral and spatial constraints. The addition of algorithmic-compression diagnostics extends this perspective beyond state entropy to measurable description length, providing a practical analogue of the system’s effective complexity. In this sense, the proposed approach translates the informational architecture emphasized by Shalizi and Crutchfield into a form suitable for empirical systems characterized by heterogeneity, nonstationarity, and feedback—hallmarks of complex adaptive social environments.
Across variables and symbolizations, results indicate that efforts and hrsncared behave as nearly memoryless processes at the dyad cadence, whereas ordinal walkability retains traces of structured but weak predictability. This suggests that short-range predictability remains limited unless (a) slower temporal baselines (e.g., hour-of-day, day-of-week, or mobility strata) are introduced, (b) embedding delays $(\tau>1)$ capture slower geometric variations, or (c) multivariate $\varepsilon$-machines jointly model walkability, mobility, and caregiving context. We adopt an ordinal symbolization scheme that stabilizes predictive information without over-merging, recognizing that detectable memory likely emerges only after conditioning or at coarser temporal scales. Overall, the results demonstrate that the joint use of $\varepsilon$-machines and compression-based diagnostics effectively separates semantic organization (predictive structure and causal memory) from syntactic simplicity (description length), providing a unified lens for examining emergence and adaptation in complex agent-based systems.
@@
Coupling $\varepsilon$-machines with compression diagnostics yields a coherent picture of where predictive information resides in the caregiving ABM. Global reconstructions furnish a baseline (often L=0L{=}0 for pooled series under coarse symbolizations), while per-dyad models reveal localized structure, especially for walkability under ordinal encodings (m=3m{=}3). Compression results corroborate this: dictionary compressors agree strongly on algorithmic redundancy, whereas normalized LZ78 highlights statistical novelty via substring diversity.
@@
Socioeconomic variables largely reflect cross-sectional heterogeneity, exhibiting near-memoryless dynamics at dyad cadence, while spatial interaction—encoded by walkability—induces bounded temporal memory and localized recurrent regimes. The contrast validates the joint diagnostic: semantic organization (predictive states and memory) is cleanly separated from syntactic simplicity (description length), clarifying how emergence manifests at different layers of the system.
@@
Symbolization is not neutral: ordinal encodings preserve temporal geometry, and hybrid Hurdle–GMM captures zero inflation. Large alphabets (m=4m{=}4) can induce effective sparsity, hiding finite memory; moderate alphabets (m=3m{=}3) with calibrated merging recover it. We recommend reporting $h_{\mu}/\log_{2}A$, $C_{\mu}$, and $E_{\mathrm{proxy}}$ alongside compression bits-per-symbol and LZ78 indices to assess stability across symbolizations.
Limitations and Robustness.
Reconstruction quality depends on symbolization, sequence length, and stationarity; short or highly entropic series can bias state estimates or collapse to trivial machines. We mitigate these issues through entropy-balanced or quantile binning, hurdle modeling for zero-inflated variables, BIC-based order selection, and complementary algorithmic proxies that retain discriminatory power when predictive structure is minimal. Possible extensions include ordinal-pattern symbolization, Bayesian smoothing of predictive distributions, and cohort pooling for short sequences.
@@
Uncertainty quantification for $(h_{\mu},C_{\mu},E_{\mathrm{proxy}})$ and compressor-based indices should be added via block bootstrap or parametric surrogates. Future extensions include multivariate $\varepsilon$-machines (jointly modeling walkability, mobility, and caregiving context), delayed embeddings ($\tau>1$) for slower rhythms, and targeted symbolizations informed by control or policy levers. Extending this framework to interactive, nonstationary, or policy-driven ABMs could help bridge informational and causal perspectives on social complexity, consistent with current efforts to quantify emergence and adaptation in empirical complex systems.

### v3

@@ 5 Case Study Outline: Caregiver–Elder Dyads
From a methodological standpoint, this demonstrates why symbolization cannot be regarded as a neutral preprocessing step. The quantile scheme was adequate for efforts, whose variability is driven primarily by cross-sectional attributes (income and elder ability) rather than by temporal feedback. In that case, the absence of causal states corresponds to genuine statistical independence. Yet for walkability and hours not cared, the same discretization distorted the data-generating process: it produced $\varepsilon$-machines that appeared memoryless despite substantive reasons to expect autocorrelation—spatial continuity in the former and dependency on caregiving load in the latter. More precisely, symbolization defines the observable coarse-graining of the underlying process. All information-theoretic quantities reported here are therefore conditional on the chosen representation, in the same sense that entropy, mutual information, and predictability are conditioned on a coarse-graining in statistical physics and information theory. Different symbolizations expose different aspects of the same underlying dynamics: coarse partitions may suppress short-range dependencies, while finer partitions may fragment trajectories and inflate entropy. The $\varepsilon$-machine reconstruction operates on the induced symbolic process and is agnostic to the physical scale at which observations are discretized.
@@
6 System-Level $\varepsilon$-machine Reconstruction and Implementation
This section applies the methodological steps summarized in Section 4. Detailed algorithmic definitions, including $\varepsilon$-machine construction, Bayesian Information Criterion (BIC) derivations, finite-block entropy estimators, and the full symbolization and discretization pipeline, are provided in Appendix A.
We report (i) a pooled, system-level reconstruction as a macro reference model and (ii) dyad-level reconstructions as micro-level complexity signatures. Implementation choices for symbolization, Markov-order selection, clustering tolerances, and entropy estimation follow the pipeline definitions in Appendix A. Throughout this section, the symbol $L$ denotes the Markov order, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm.
6.1 System-Level (Pooled) Reconstruction as a Macro Reference
Let $\mathbf{s}^{\mathrm{pool}}=(s_{1},\dots,s_{n})$ denote a pooled symbolic sequence obtained by aggregating the selected variable across dyads and time after applying a fixed symbolization scheme (Appendix A). We reconstruct a single $\varepsilon$-machine from $\mathbf{s}^{\mathrm{pool}}$ and report the corresponding $(h_{\mu},C_{\mu},E_{\mathrm{proxy}})$ as a compact summary of the system-wide informational regime under that observational resolution.
The Markov order $L^{\ast}$ is selected by minimizing $\mathrm{BIC}(L)$ over $L\in\{0,\dots,L_{\max}\}$, and causal states are obtained by clustering empirical predictive distributions using the $\ell_{1}$ norm (see Appendix A). When $L^{\ast}=0$, the pooled process behaves approximately memoryless under the chosen symbolization. We interpret this case as a macro baseline: it indicates that, at the pooled level and given the current discretization, predictive structure is not detectable beyond zero-order dependence, motivating the micro-level analysis below.
6.2 Dyad-Level Reconstruction and Multi-Variable Feature Extraction
For each dyad $i$ and each variable $v\in\{\texttt{efforts},\texttt{wkb},\texttt{hrsncared},\texttt{overwhelmed}\}$, we construct a symbolic sequence $\mathbf{s}^{(i,v)}$ using the same family of discretization rules (Appendix A) and reconstruct an $\varepsilon$-machine for $\mathbf{s}^{(i,v)}$. From each reconstructed model we extract the entropy rate, statistical complexity, and predictive information proxy, yielding the dyad- and variable-specific signature
$\big(h_{\mu}^{(i,v)},\,C_{\mu}^{(i,v)},\,E_{\mathrm{proxy}}^{(i,v)}\big).$
These quantities enable distributional analysis across dyads, stratified comparisons by exogenous attributes (e.g., stage, disability), and unsupervised clustering of behavioral regimes.
6.3 Algorithmic Complexity Proxies from Lossless Compression
Because $\varepsilon$-machine reconstructions may collapse to trivial structures for short or weakly dependent sequences, we complement the causal-state quantities with algorithm-agnostic proxies for descriptive regularity. For each symbolic sequence $\mathbf{s}^{(i,v)}$, we compute: (i) normalized LZ78 complexity; and (ii) bits-per-symbol estimates from multiple lossless codecs (LZMA, BZ2, GZIP), defined as
$C_{\mathrm{comp}}=\frac{8\,S_{\mathrm{comp}}}{n},$ (8)
where $S_{\mathrm{comp}}$ is the compressed size in bytes and $n$ is the sequence length. Using multiple codecs provides a practical multi-scale view of algorithmic regularity, since LZMA tends to capture longer-range structure, BZ2 captures medium-range regularities, and GZIP is dominated by short-range redundancy (Appendix A).
6.4 Outputs, Reporting Scope, and Reproducible Artifacts
All reported metrics are produced at two hierarchical scopes: (i) the pooled system level and (ii) the dyad level, with optional aggregation to meso-level strata. For dissemination and auditability, we provide: (a) distributions and heatmaps of $\{h_{\mu},C_{\mu},E_{\mathrm{proxy}},\mathrm{LZ78},C_{\mathrm{comp}}\}$ by variable and dyad; (b) a concise diagram of the pooled $\varepsilon$-machine as a macro reference; and (c) tabular summaries of $L^{\ast}$ and model size ($|S|$) for each reconstruction. All algorithmic definitions and implementation details needed to reproduce these outputs are given in Appendix A.
@@ 8 Discussion, Limitations, and Conclusions
This study applies tools from computational mechanics to the analysis of agent-based model (ABM) outputs, with the specific goal of diagnosing temporal organization and predictive structure in simulated social dynamics. While $\varepsilon$-machine reconstruction has been developed primarily in the context of symbolic and nonlinear dynamical systems [10, 11, 12], its use in applied ABM analysis remains limited. The present work demonstrates how these methods can be operationalized within a data-driven simulation workflow, without relying on theoretical toy processes or analytically specified sources.
Methodologically, the framework integrates symbolic dynamics, causal-state reconstruction, and algorithmic compression measures to characterize the informational organization of ABM-generated time series. Each component targets a distinct aspect of structure: $\varepsilon$-machines identify predictive equivalence classes of histories; entropy- and complexity-based quantities summarize information generation and storage; and compression-based indices quantify description length and redundancy. Taken together, these diagnostics provide a reproducible and internally consistent description of temporal regularity across variables, agents, and observational resolutions.
A central distinction clarified by the analysis is that between predictive structure and descriptive regularity. Following computational mechanics, intrinsic computation concerns how past observations constrain future behavior through causal states [11, 12]. Compression-based measures, by contrast, estimate the compactness of a symbolic description without implying causal predictability. The joint application of these tools therefore serves a diagnostic purpose: agreement between them indicates robust structure, whereas divergence highlights scale-, representation-, or variable-specific effects.
Empirically, the results indicate that several variables produced by the caregiving ABM—notably efforts and hours not cared—exhibit near-memoryless behavior at the dyad time scale under the symbolizations considered. In contrast, walkability, when encoded using ordinal patterns, retains weak but detectable predictive structure. This suggests that temporal organization in the model is variable-specific and closely tied to spatial and interaction-driven processes rather than to purely cross-sectional attributes. Importantly, the absence of detectable memory at a given resolution does not imply an absence of structure in the underlying dynamics, but rather reflects the conditional nature of inference under a chosen coarse-graining.
These findings underscore the role of symbolization as an explicit modeling assumption. Symbolization defines the observable coarse-graining of the process, and all reported information-theoretic quantities are conditional on this representation, in the same sense that entropy and mutual information depend on the chosen partition of state space in information theory [Shannon1948, 13]. Ordinal encodings preserve local temporal geometry, while hybrid schemes such as hurdle–mixture discretizations address structural zeros. No symbolization is privileged as “correct”; rather, different encodings expose different aspects of the same simulated dynamics.
@@
At the system level, pooled reconstructions often select $L^{\ast}=0$ under coarse symbolizations, yielding a macro-level baseline that is effectively memoryless. At the dyad level, however, $\varepsilon$-machine reconstructions reveal heterogeneity across variables and agents, with nontrivial predictive structure emerging primarily for spatially mediated quantities under ordinal symbolizations. Compression-based diagnostics corroborate these patterns by distinguishing between redundancy and novelty across sequences, even when causal state structure is minimal.
@@
The contrast between socioeconomic and spatial variables suggests that temporal organization in the model arises primarily from interaction with the simulated environment rather than from static agent attributes. This separation between predictive memory and descriptive simplicity clarifies how different forms of structure coexist in ABM outputs and how they can be disentangled using complementary diagnostics.
@@
The analysis highlights the importance of reporting multiple, non-redundant complexity measures. We recommend presenting entropy-rate–normalized quantities, statistical complexity, predictive-information proxies, and compression-based bits-per-symbol jointly, to assess robustness across symbolizations and scales. Reliance on a single metric risks conflating descriptive redundancy with predictive organization.
Limitations and Scope
The results presented here are subject to several important limitations. First, all analyses are simulation-specific. The symbolic sequences analyzed are generated entirely by an agent-based model and do not constitute empirical observations. Accordingly, the reported measures characterize the informational organization of the simulated dynamics under controlled assumptions, rather than explaining or predicting real-world caregiving behavior.
Second, all information-theoretic quantities are representation-dependent. Symbolization defines the observable coarse-graining of the underlying process, and entropy rates, statistical complexity, and predictive information are conditional on this choice, as is standard in information theory and statistical physics [17, 13]. Alternative discretizations may reveal different aspects of the same underlying dynamics.
Third, the framework is explicitly diagnostic rather than explanatory. $\varepsilon$-machine reconstructions identify predictive equivalence classes of histories and quantify temporal organization, but they do not identify causal mechanisms in an interventionist sense, nor do they constitute generative models of agent decision-making [11, 12]. The reported measures should therefore be interpreted as descriptive summaries suited to comparison, stratification, and robustness analysis.
@@
Future work should incorporate uncertainty quantification for entropy, complexity, and compression-based measures, for example via block bootstrap or surrogate-based methods. Methodological extensions include multivariate $\varepsilon$-machines, delayed embeddings to capture slower rhythms, and symbolizations informed by policy-relevant thresholds. Applying the framework to nonstationary or intervention-driven ABMs may further clarify how informational structure responds to structural change in simulated social systems.
@@ Acknowledgments
Appendix A Algorithmic Definitions and Implementation Details
This appendix provides the full algorithmic definitions referenced in Sections 47. It collects reconstruction, model-selection, and entropy-estimation details to avoid redundancy in the main text.
Simulation Configuration
Table 6 summarizes the core configuration parameters of the agent-based simulation used to generate all time series analyzed in this study. These parameters define the scale and temporal resolution of the simulated system and are reported to facilitate replication and contextual interpretation of the results. All analyses are conditional on this simulation design.
Table 6: Core simulation configuration parameters.
Parameter Value / Description
Simulation type Agent-based model of caregiver–elder dyads
Unit of analysis Caregiver–elder dyad
Number of dyads Fixed per simulation run (order of $10^{2}$)
Time step Discrete simulation tick (uniform temporal resolution)
Simulation horizon Fixed number of ticks per run (order of $10^{3}$)
Recorded variables Caregiving effort (efforts), walkability index (wkb), hours not cared (hrsncared), overwhelmed status (binary)
Variable nature Mixed continuous, zero-inflated, and binary variables
Spatial structure Explicit simulated physical environment with distances and service locations
Initialization Socioeconomic attributes fixed at initialization; behavioral and accessibility variables evolve over time
Data origin Fully simulated; no empirical or observational data used
Compression-Based Complexity Proxies: Implementation Details
Compression-based measures are employed in this study as algorithm-agnostic proxies for descriptive regularity and algorithmic complexity in symbolic sequences generated by the agent-based model. These measures complement $\varepsilon$-machine–based quantities by quantifying redundancy and compressibility without making assumptions about predictive structure or causal state representations.
All compression algorithms are applied using standard library implementations with default parameter settings. [^1] Default compression settings are used intentionally. Parameter tuning can improve compression ratios for particular sequence lengths or symbol distributions, but it also introduces algorithm- and data-dependent bias. For diagnostic studies aimed at comparative analysis rather than optimal encoding, standardized defaults ensure neutrality, reproducibility, and consistency across compression algorithms. This choice is intentional: while parameter tuning can improve compression performance for specific sequence lengths or symbol distributions, it also introduces algorithm- and data-dependent bias. In a diagnostic and comparative context—where the objective is to assess relative structure across variables, dyads, and representations rather than to achieve optimal encoding—the use of standardized defaults preserves methodological neutrality, reproducibility, and comparability across algorithms.
For each symbolic sequence of length $n$, the compressed output size $S_{\mathrm{comp}}$ (in bytes) is converted to a normalized measure of bits per symbol,
$C_{\mathrm{comp}}=\frac{8\,S_{\mathrm{comp}}}{n},$ (9)
allowing direct comparison with entropy-rate estimates expressed in bits per symbol. Compression is performed on raw byte streams representing the symbolic sequences, without metadata or container-specific headers.
Table 7: Compression-based complexity proxies: algorithms and parameter settings.
Algorithm Implementation Key parameters Values used and rationale
LZMA (Lempel–Ziv–Markov) Python lzma (XZ Utils reference implementation) Compression preset; dictionary size; container format Default preset (preset = 6); dictionary size determined by preset (library default); FORMAT_XZ. Defaults balance compression strength and computational cost and allow detection of long-range dependencies without algorithm-specific tuning.
BZ2 (Burrows–Wheeler transform) Python bz2 Block size; entropy coding Default block size (blocksize = 9, 900 kB); Huffman coding enabled by default. Captures medium-range regularities with moderate computational cost.
GZIP (DEFLATE: LZ77 + Huffman) Python gzip Compression level; sliding window Default maximum compression level (compresslevel = 9); sliding window fixed by DEFLATE standard (32 kB). Emphasizes short-range redundancy and fast execution.
All compressors Input format; tuning policy; output metric Symbol sequences serialized as raw byte streams without metadata. No parameter tuning performed beyond library defaults to preserve comparability across variables and dyads. Compressed size $S_{\mathrm{comp}}$ normalized as bits per symbol: $C_{\mathrm{comp}}=\frac{8\,S_{\mathrm{comp}}}{n},$ where $n$ is the sequence length.
A.1 System-Level $\varepsilon$-machine Reconstruction: Definitions
$h_{\mu}=H[X_{t}\mid X_{:t}],$ (11)
$C_{\mu}=H[\mathcal{S}],$ (12)
$E=I[X_{:t};\,X_{t:\infty}],$ (13)
Accurate estimation of these quantities, however, critically depends on how continuous simulation outputs are represented as discrete symbolic sequences. Symbolization—the transformation of numeric or multidimensional trajectories into finite alphabets—acts as the analytical lens through which temporal organization becomes observable. An appropriate scheme determines which variations are treated as equivalent and which are preserved as meaningful distinctions. Poorly chosen symbolizations can obscure dependencies by merging distinct states or, conversely, fabricate structure by discretizing noise into artificial categories.
In agent-based simulations, this step is particularly delicate because variables differ in scale and generative mechanism. Spatial indicators such as walkability exhibit smooth continuity and feedback, requiring ordinal or rank-based encodings that preserve relative change, whereas socioeconomic measures like caregiver effort or hours not cared vary primarily across agents and are better discretized using quantile or hurdle-mixture schemes. Symbolization thus defines the effective alphabet and the grain of observation, balancing detail and parsimony. Finer encodings may inflate entropy and complexity through over-partitioning, while coarser ones suppress correlations and underestimate structure.
Beyond its technical role, symbolization is a modeling decision: it specifies the mapping between the simulated micro-dynamics and the informational architecture that $\varepsilon$-machines can reveal. From an information-theoretic perspective, this dependence on representation is not a methodological flaw but a standard conditioning assumption. Entropy and related quantities are defined relative to an observation alphabet, and coarse-graining alters the measurable information content of a process without changing its underlying generative mechanism. This is analogous to the role of macrostates in statistical physics, where entropy depends on the chosen partition of phase space rather than on microscopic dynamics [cover2006elements, 17].
A.2 $\varepsilon$-Machine Reconstruction and Entropy Estimation Pipeline
A.2.1 $\varepsilon$-Machine Reconstruction
For a given Markov order $L$, we denote by $s_{t-L:t-1}$ the length-$L$ history of symbols immediately preceding time $t$.
Let $N(h,a)$ denote the number of occurrences of symbol $a\in\{0,\dots,A{-}1\}$ following a history $h=s_{t-L:t-1}$, and let $N(h)=\sum_{a}N(h,a)$ denote the total number of occurrences of history $h$. For each candidate order $L$, the empirical next-symbol probabilities are then defined as:
$\hat{P}(a\mid h)=\frac{N(h,a)}{N(h)},$ (14)
The corresponding log-likelihood and Bayesian Information Criterion are given by:
$\log\mathcal{L}_{L}=\sum_{h}\sum_{a}N(h,a)\,\log\hat{P}(a\mid h),$ (15)
$\mathrm{BIC}(L)=-2\log\mathcal{L}_{L}+k_{L}\log N_{\mathrm{eff}},$ (16)
All subsequent reconstruction steps, including history clustering, transition construction, and information-theoretic estimation, are performed using the selected Markov order $L^{\ast}$ unless explicitly stated otherwise.
A.2.2 Predictive Distributions and History Clustering
For the selected order, predictive conditional distributions $P(a_{t+1}\mid a_{t-L:t})$ are estimated for all histories observed with sufficient frequency. Histories with statistically similar predictive distributions (as measured by their $\ell_{1}$ norm) are clustered to form equivalence classes representing causal states.
For each observed history $h$, we estimate:
$\hat{P}(a\mid h)=\frac{N(h,a)}{N(h)},\qquad a\in\{0,\dots,A{-}1\},$ (17)
where $N(h,a)$ is the joint count of occurrences of history $h$ followed by symbol $a$, and $N(h)$ is the total count of $h$. Histories observed fewer than a threshold count are discarded to suppress noise. Two histories $h$ and $h^{\prime}$ belong to the same causal state if their $\ell_{1}$ norm satisfies
$\hat{P}_{\mathrm{emit}}(a\mid s)=\frac{1}{|C_{s}|}\sum_{h\in C_{s}}\hat{P}(a\mid h),$ (19)
defaulting to a uniform vector if no histories are present. The distribution $\hat{P}_{\mathrm{emit}}(\cdot\mid s)$ represents the empirical symbol emission frequencies associated with causal state $s$ and does not constitute a transition probability between states.
A.2.3 Unifilar Transitions and Stationary Distribution
A unifilar transition structure is a state–transition topology used in computational mechanics and predictive-state models (such as $\varepsilon$-machines), where each state–symbol pair uniquely determines the next state. In empirical reconstructions from finite data, unifilarity is enforced approximately by constructing a deterministic transition function via a majority rule over observed history shifts. The transition function $T(s_{i},a)=s_{j}$ is deterministic: it specifies the unique next causal state reached from state $s_{i}$ upon emission of symbol $a$. Transition probabilities are induced indirectly via the emission distributions $\hat{P}_{\mathrm{emit}}(\cdot\mid s_{i})$ and the stationary distribution $\pi$, and are not associated with $T$ itself.
$h^{\prime}=\mathrm{shift}_{L}(h,a),$ (20)
where $\mathrm{shift}_{L}(\cdot,\cdot)$ appends symbol $a$ to $h$ and discards the oldest symbol to preserve history length $L$. The next state $s^{\prime}$ is then chosen as the most frequently observed causal state reached by these shifted histories. This procedure yields a deterministic mapping $T(s,a)=s^{\prime}$ for each state–symbol pair.
The corresponding state-transition matrix $M$ is defined as
$M_{ij}=\sum_{a}\hat{P}_{\mathrm{emit}}(a\mid s_{i})\,\mathbf{1}\!\{T(s_{i},a)=s_{j}\},$ (21)
and is normalized row-wise to ensure stochasticity. The stationary distribution $\pi$ is obtained as the fixed point of the Markov chain,
$\pi=\pi M,\qquad\pi^{(t+1)}=\pi^{(t)}M,\qquad\|\pi^{(t+1)}-\pi^{(t)}\|_{1}<\mathrm{tol},$ (22)
where convergence is assessed using the $\ell_{1}$ norm.
A.2.4 Optional State Merging by Jensen–Shannon Divergence
To regularize over-fragmented models, states with similar emission distributions are merged when
$\mathrm{JS}\!\big(\hat{P}_{\mathrm{emit}}(\cdot\mid s),\,\hat{P}_{\mathrm{emit}}(\cdot\mid s^{\prime})\big)\leq\tau_{\mathrm{JS}},$ (23)
where $\mathrm{JS}(P,Q)$ denotes the Jensen–Shannon divergence and $\tau_{\mathrm{JS}}$ controls the merging threshold. The divergence is defined as
and after merging, the stationary distribution $\pi$ is recomputed.
A.2.5 Information-Theoretic Quantities
Entropy rate.
$h_{\mu}=\sum_{s}\pi_{s}\,H\!\big(\hat{P}_{\mathrm{emit}}(\cdot\mid s)\big),\qquad H(p)=-\sum_{i}p_{i}\log_{2}p_{i},$ (25)
quantifying the average unpredictability of symbol generation under the stationary distribution of causal states.
Statistical complexity.
$C_{\mu}=H(\pi)=-\sum_{s}\pi_{s}\log_{2}\pi_{s},$ (26)
measuring the amount of information stored in the stationary causal-state distribution.
Predictive information proxy.
Exact computation of the excess entropy $E=I[X_{:t};X_{t:\infty}]$ is generally intractable for finite reconstructed models. We therefore report a one-step predictive information proxy based on the reconstructed $\varepsilon$-machine. Let
$\hat{P}_{\mathrm{next}}=\sum_{s}\pi_{s}\hat{P}_{\mathrm{emit}}(\cdot\mid s),\qquad H_{\mathrm{next}}=H(\hat{P}_{\mathrm{next}}),$ (27)
then
$E_{\mathrm{proxy}}=\max\!\big(H_{\mathrm{next}}-h_{\mu},\,0\big),$ (28)
which approximates the amount of predictive structure captured at the one-step horizon. By construction, $E_{\mathrm{proxy}}$ provides a lower-bound proxy for the true excess entropy and is reported alongside block-based estimates below.
Finite-block entropy estimates and excess-entropy approximations are reported only as auxiliary diagnostics and are provided in Appendix A for completeness. They are not used for inference or model selection in the main analysis.
A.2.6 Visualization and Data Export
  1. (i)
    a time-series plot of the input symbols;
  2. (ii)
    a bar plot of stationary causal-state probabilities;
  3. (iii)
  4. (iv)
  5. (v)
    a table of BIC scores across candidate Markov orders.
The pipeline returns the selected Markov order $L^{\ast}$, the number of causal states $|S|$, the entropy rate $h_{\mu}$, statistical complexity $C_{\mu}$, the predictive information proxy $E_{\mathrm{proxy}}$, block-based estimates $\big(h_{\mu}^{(\mathrm{block})},E^{(\mathrm{block})}\big)$, and file paths for all generated visual and tabular artifacts.
A.3 Symbolization and Discretization Pipeline
Let the numeric series be denoted by
$\mathbf{x}=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}.$ (29)
A binning scheme with $k$ bins is defined by the ordered set of edges
$\mathbf{e}=(e_{0},\dots,e_{k}),\qquad e_{0}<\cdots<e_{k}.$ (30)
A.3.1 Discretization Methods
[…]
Here, the parameter $m$ denotes the local embedding or pattern order used for symbolization and is conceptually distinct from the Markov order $L$ employed later in $\varepsilon$-machine reconstruction.
Fallback mechanisms ensure stability: Jenks $\rightarrow$ K-Means $\rightarrow$ quantiles for class breaks, and default clipping to small perturbations when numerical degeneracy occurs. These fallback procedures are purely algorithmic safeguards to ensure numerical stability and do not introduce additional modeling assumptions.
1. Quantile Bins.
$e_{j}=Q\!\left(\frac{j}{k}\right),\qquad j=0,\dots,k.$ (32)
2. Fixed or Policy-Fixed Edges.
3. Winsorized Equal-Width.
Using parameters $p_{\mathrm{low}}$ and $p_{\mathrm{high}}$, compute
$\ell=Q(p_{\mathrm{low}}),\qquad h=Q(p_{\mathrm{high}}),$ (33)
then clip as in Section A.3.2. Edges are constructed as
$e_{j}=\ell+\frac{j}{k}(h-\ell),\qquad j=0,\dots,k.$ (34)
Alphabet size: $A=k$.
4. Jenks Natural Breaks.
Use edges $\mathbf{e}^{\star}$ obtained from the Jenks optimization in Section A.3.2. Assign labels $\ell_{i}=\ell(x_{i};\mathbf{e}^{\star})$ and set alphabet size $A=k$.
5. Hurdle + Gaussian Mixture Model.
and shift labels by one to preserve the zero bin:
Alphabet size: $A=\hat{m}+1$.
6. Ordinal Patterns.
For an order parameter $m$, construct sliding windows
$\mathbf{w}_{t}=(x_{t-m+1},\dots,x_{t}).$ (39)
7. Lagged Vector Quantization.
For lag order $m$, build lagged vectors
$\mathbf{y}_{t}=(x_{t-m+1},\dots,x_{t})^{\top},\qquad t=m,\dots,n.$ (41)
Optionally standardize as
$\mathbf{z}_{t}=(\mathbf{y}_{t}-\hat{\mu})/\hat{\sigma}$, (42)
(elementwise). Apply $k_{\mathrm{VQ}}$-means clustering to minimize
$\sum_{t=m}^{n}\min_{j}\|\mathbf{z}_{t}-\mathbf{c}_{j}\|_{2}^{2}.$ (43)
Assign each vector to its nearest centroid:
$\ell_{t}=\arg\min_{j}\|\mathbf{z}_{t}-\mathbf{c}_{j}\|_{2}^{2}.$ (44)
A.3.2 Binning, Digitization, and Label Formation
Digitization.
Given edges $\mathbf{e}$, integer labels are computed as
Equal-Width Edges.
$e_{j}=\ell+\frac{j}{k}(h-\ell),\qquad j=0,\dots,k.$ (46)
Jenks Natural Breaks.
Jenks natural breaks [15] minimize the within-class sum of squares:
Winsorization.
$\ell=Q(p_{\text{low}}),\qquad h=Q(p_{\text{high}}),$ (49)
$x_{i}^{\text{win}}=\min\{\max\{x_{i},\ell\},\,h\}.$ (50)
The function returns $(\mathbf{x}^{\text{win}},\ell,h)$.
Appendix B Finite-Block Entropy and Excess Entropy Diagnostics
This appendix reports finite-block entropy estimates as a diagnostic consistency check against $\varepsilon$-machine–based quantities.
An independent estimate of the entropy rate and excess entropy is obtained via finite-block entropies up to block length $L_{\max}$. For each block length $L$, the empirical block entropy $H_{L}$ is estimated using a plug-in estimator with mild additive smoothing. Asymptotic values are approximated as
$h_{\mu}^{(\mathrm{block})}\approx H_{L}-H_{L-1},\qquad E^{(\mathrm{block})}\approx H_{L}-L\,h_{\mu}^{(\mathrm{block})},$ (51)
which follow standard finite-difference approximations for stationary symbolic processes.
$\hat{P}(b)=\frac{N(b)+1}{N+|\mathcal{B}^{L}|},$ (52)
and the corresponding block entropy is computed as
$H_{L}=-\sum_{b}\hat{P}(b)\log_{2}\hat{P}(b).$ (53)
To reduce finite-sample noise, block-based estimates are refined by averaging the final three increments of $H_{L}-H_{L-1}$: